OpenStudy (x3_drummerchick):

how do i start to solve this? will give medals! Find the first 4 partial sums of the sequence, then find the value for S(sub n): a(sub n) = 2/(3^n)

OpenStudy (x3_drummerchick):



OpenStudy (x3_drummerchick):

@Nnesha

OpenStudy (anonymous):

A partial sum \(S_k\) for a sequence \(a_n\) is given by \[\sum_{n=1}^ka_n=a_1+a_2+a_3+\cdots+a_k\] We call \(S_k\) the \(k\)th partial sum for the sequence \(a_n\). The first partial sum, for example, is simply \[\sum_{n=1}^1a_n=a_1=\frac{2}{3^1}=\frac{2}{3}\] Make sense?